OpenStudy (anonymous):

simplify 3 squared root of 2x + x square root of of 8x - 5 square root of 18x

OpenStudy (hoblos):

3√2x + x√8x -5√18x = 3√2x + x√(4*2x) -5√(9*2x) = 3√2x + 2x√2x -(3*5)√2x = ( 3+2x -15)√2x = ( 2x-12)√2x

OpenStudy (anonymous):

thanks!